OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x + 9

OpenStudy (demonchild99):

y =mx+b so just plug in the numbers in the correct spots and there ya go

OpenStudy (demonchild99):

what do you think saifoo?

OpenStudy (demonchild99):

@saifoo.khan

OpenStudy (saifoo.khan):

Step 1: Find the gradient. Step 2: Use the formula: (y-y1) = m(x-x1)